# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v1

@@ I Introduction
Figure 2: Illustration of our proposed framework. With Optimizer 1, we simultaneously train the supervised Syn-Real depth estimation and the Color Net to predict colors from depths in real world. With Optimizer 2, we train real-world unsupervised depth estimation with sky regularization loss.
@@ III-A1 Transform depth to specific colors
And utilizing the input of the Depth Net $I^{\mathcal{X}}$ for supervision:
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.III, Tab.IV, and Tab.I). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.VIII), perform comparisons at higher resolutions (Tab.V), and benchmark against the lightest and most effective methods (Tab.IX). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 11, Fig. 14, and Fig. 15).
@@ IV-C Monocular Depth Estimation Performance
The evaluation results are summarized in Tab.III and Tab.IV with Kitti Eigen Split[47] and Tab.I with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.11 and Fig.14. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[34, 33] with larger params and more complex structures(Tab.IX).
@@ IV-D Ablation Study
Method Supervision Datasets Size Error Metric Accuracy Metric
@@
Monodepth2[8] m K 192x640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
CADepth[24] m K 192x640 0.105 0.769 4.535 0.181 0.892 0.964 0.983
DIFFNet[26] m K 192x640 0.102 0.749 4.445 0.179 0.897 0.965 0.983
MonoFormer[62] m K 192x640 0.106 0.839 4.627 0.183 0.889 0.962 0.983
Lite-Mono[34] m K 192x640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
MonoViT[33] m K 192x640 0.099 0.708 4.372 0.175 0.900 0.967 0.984
Back2Color m K+SHIFT 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
DIFFNet[26] m K 320x1024 0.097 0.722 4.345 0.174 0.907 0.967 0.984
Lite-[34] m K 320x1024 0.097 0.710 4.309 0.174 0.905 0.967 0.984
MonoViT[33] m K 320x1024 0.096 0.714 4.292 0.172 0.908 0.968 0.984
Back2Color m K+vKitti 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
@@
Method Supervision Datasets Size Error Metric Accuracy Metric
@@
Monodepth2[8] ms K 192x640 0.106 0.818 4.750 0.196 0.874 0.957 0.979
HR-Depth[25] ms K 192x640 0.107 0.785 4.612 0.185 0.887 0.962 0.982
CADepth[24] ms K 192x640 0.102 0.752 4.504 0.181 0.894 0.964 0.983
DIFFNet[26] ms K 192x640 0.101 0.749 4.445 0.179 0.898 0.965 0.983
MonoViT[33] ms K 192x640 0.098 0.683 4.333 0.174 0.904 0.967 0.984
Back2Color ms K+SHIFT 192x640 0.092 0.676 4.238 0.171 0.909 0.969 0.984
Monodepth2[8] ms K 320x1024 0.106 0.806 4.630 0.193 0.876 0.958 0.980
HR-Depth[25] ms K 320x1024 0.101 0.716 4.395 0.179 0.894 0.966 0.983
CADepth[24] ms K 320x1024 0.096 0.694 4.264 0.173 0.908 0.968 0.984
MonoViT[33] ms K 320x1024 0.093 0.671 4.202 0.169 0.912 0.969 0.984
Back2Color ms K+SHIFT 320x1024 0.089 0.651 4.158 0.167 0.916 0.970 0.985
@@
Method Supervision Error Metric Accuracy Metric
@@
Back2Color DIFFNet 0.097 0.690 4.295 0.175 0.901 0.966 0.984
@@
Back2Color VAN2 0.092 0.676 4.238 0.171 0.909 0.969 0.984
TABLE VII: Ablation experiment on the number of dataset samples, the total number used for unsupervised training in the KITTI dataset is 39,810. We replaced 20% and 50% of it with synthetic datasets respectively.
Syn data $N_{Real}$ $N_{Syn}$ Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
vKitti2[10] 31848 7962 0.090 0.675 4.143 0.168 0.917 0.970 0.985
vKitti2[10] 19905 19905 0.091 0.684 4.192 0.169 0.913 0.969 0.985
SHIFT[45] 31848 7962 0.090 0.684 4.158 0.168 0.916 0.970 0.984
SHIFT[45] 19905 19905 0.090 0.680 4.144 0.167 0.915 0.970 0.985
@@
Additionally, we used various synthetic datasets such as vKitti, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Fig.VIII. The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method.
Finally, to demonstrate that the improvement in our Back2Color method is not due to the increased training samples(Synthetic Datasets), we conducted experiments using 20% and 50% synthetic samples while ensuring the total number of dataset samples remains the same as the KITTI dataset (39,810), as shown in Tab.VII. The results did not decline, proving that the enhancement of our method is not due to the increase in training samples.
TABLE VIII: Ablation study on different synthetic datasets, using temporal and stereo sequences(ms) for supervision
Size Datasets Error Metric Accuracy Metric
abs_rel sq_rel rmse rmse_log a1 a2 a3
192x640 K+vKitti 0.094 0.694 4.296 0.173 0.907 0.968 0.984
192x640 K+SHIFT 0.092 0.676 4.238 0.171 0.909 0.969 0.984
192x640 K+Urban 0.093 0.644 4.198 0.172 0.907 0.968 0.984
320x1024 K+vKitti 0.090 0.677 4.173 0.169 0.914 0.969 0.984
320x1024 K+SHIFT 0.089 0.651 4.158 0.167 0.916 0.970 0.985
320x1024 K+Urban 0.090 0.654 4.084 0.167 0.915 0.970 0.985
@@ IV-E Visualization Analysis
TABLE IX: Ablation study for models with params and computational complexities(GFLOPs).
@@
MonoViT[33] small 320x1024 33.63 159.14 0.093 0.671 4.202 0.169 0.912 0.969 0.985
Lite-mono[34] 8M 320x1024 8.76 29.88 0.097 0.710 4.309 0.174 0.905 0.967 0.984
Back2Color VAN0 320x1024 5.17 17.64 0.097 0.749 4.357 0.174 0.905 0.968 0.984
Back2Color VAN1 320x1024 18.61 51.48 0.091 0.676 4.177 0.170 0.913 0.969 0.984
Back2Color VAN2 320x1024 31.32 76.01 0.089 0.651 4.158 0.167 0.916 0.970 0.985
Back2Color VAN3 320x1024 49.5 114.63 0.088 0.661 4.114 0.167 0.917 0.970 0.984
Back2Color VAN0 384x1280 5.17 17.64 0.096 0.712 4.269 0.173 0.908 0.969 0.984
Back2Color VAN1 384x1280 18.61 51.48 0.094 0.703 4.187 0.171 0.911 0.969 0.984
Back2Color VAN2 384x1280 31.32 76.01 0.088 0.613 4.049 0.165 0.918 0.971 0.985
Back2Color VAN3 384x1280 49.5 114.63 0.088 0.668 4.103 0.167 0.918 0.970 0.984
@@ IV-F3 Can the VAN-based encoder replace the ResNet structure in the Pose encoder
so far, we have not observed any new research that can replace the ResNet-based PoseEncoder, whether using CNN-based or Transformer-based methods. We have attempted some alternatives, but none have improved depth estimation performance; in fact, many have resulted in performance degradation. In this study, we also tried using a VAN-based model to replace the ResNet-based PoseEncoder, but still did not observe any significant improvement. We believe that the pose estimation task is a small parameter regression task that is relatively simple compared to depth prediction and does not require a powerful network.

### v2

@@ I Introduction
Figure 2: Illustration of our framework. Using Optimizer 1, we train supervised Syn-Real depth estimation and the Color Net to predict real colors from depths. First, ①we fix the Depth Net and Color Net to infer real and synthetic predicted depths, using the synthetic depth to predict Syn-to-Real Color. Then, ②we apply Syn-Real Cutmix to create new syn-real colors, ③optimizing the Depth Net with synthetic groundtruth and the Color Net on real dataset. With Optimizer 2, ④we train real-world unsupervised depth estimation with sky regularization loss and auto-learning uncertainty. The dotted line indicates inference only.
@@ III-A1 Transform depth to specific colors
And utilizing the input of the Depth Net $I^{\mathcal{X}}$ for Sup:
@@ III-D2 Optimization Strategy
Figure 11: The MixSyn dataset contains more samples compared to the real-world unsupervised datasets, necessitating the division of the MixSyn dataset into multiple epochs. While the MixSyn dataset is trained through once, the real-world datasets may undergo multiple training passes.
@@ IV Experiments
In this section, we initially establish the superiority of our Back2Color by comparing it with other state-of-the-art (SoTA) approaches on the KITTI Eigen split and Cityscapes depth estimation benchmark (Tab.III, Tab.IV, and Tab.I). Following this, we conduct ablation experiments to analyze the effects of the proposed methods (Tab.VII and Tab.VI). Additionally, we compare our approach with other methods that utilize synthetic datasets (Tab.II), perform comparisons at higher resolutions (Tab.V), and benchmark against the lightest and most effective methods (Tab.VIII). At the same time, we present some visual results of our method on the KITTI and Cityscapes datasets (Fig. 12, Fig. 15, and Fig. 16).
@@ IV-A Datasets
To leverage the variety of multiple synthetic datasets, we normalized and shuffled the specifications (size, depth resolution, etc.) of the vKitti2, SHIFT, and UrbanSyn datasets to create a unified dataset called MixSyn. The MixSyn dataset contains significantly more samples compared to KITTI and Cityscapes. Therefore, during the training process, MixSyn is trained once, while KITTI or Cityscapes is trained for multiple epochs, as shown in Fig.11. We demonstrate the advantages of using MixSyn in Tab. VII.
@@ IV-C Monocular Depth Estimation Performance
Our recommended experimental setup is to use the VAN2 model (balancing efficiency and performance), utilize the MixSyn synthetic dataset, and employ spatiotemporal supervision (ms). If there are no annotations, the results should be based on this setup. The evaluation results are summarized in Tab.III and Tab.IV with Kitti Eigen Split[47] and Tab.I with Cityscapes benchmark, followed by an illustration of their performance qualitatively in Fig.12 and Fig.15. Our method outperforms all existing SoTA unsupervised approaches, including some based on transformer[34, 33] with larger params and more complex structures(Tab.VIII).
@@ IV-D Ablation Study
Method Sup. Size Error Metric Accuracy Metric
@@
Monodepth2[8] m 192x640 0.115 0.903 4.863 0.193 0.877 0.959 0.981
CADepth[24] m 192x640 0.105 0.769 4.535 0.181 0.892 0.964 0.983
DIFFNet[26] m 192x640 0.102 0.749 4.445 0.179 0.897 0.965 0.983
MonoFormer[62] m 192x640 0.106 0.839 4.627 0.183 0.889 0.962 0.983
Lite-Mono[34] m 192x640 0.101 0.729 4.454 0.178 0.897 0.965 0.983
MonoViT[33] m 192x640 0.099 0.708 4.372 0.175 0.900 0.967 0.984
Back2Color m 192x640 0.096 0.708 4.307 0.173 0.905 0.968 0.984
DIFFNet[26] m 320x1024 0.097 0.722 4.345 0.174 0.907 0.967 0.984
Lite-[34] m 320x1024 0.097 0.710 4.309 0.174 0.905 0.967 0.984
MonoViT[33] m 320x1024 0.096 0.714 4.292 0.172 0.908 0.968 0.984
Back2Color m 320x1024 0.093 0.687 4.171 0.168 0.912 0.970 0.985
@@
Method Sup. Size Error Metric Accuracy Metric
@@
Monodepth2[8] ms 192x640 0.106 0.818 4.750 0.196 0.874 0.957 0.979
HR-Depth[25] ms 192x640 0.107 0.785 4.612 0.185 0.887 0.962 0.982
CADepth[24] ms 192x640 0.102 0.752 4.504 0.181 0.894 0.964 0.983
DIFFNet[26] ms 192x640 0.101 0.749 4.445 0.179 0.898 0.965 0.983
MonoViT[33] ms 192x640 0.098 0.683 4.333 0.174 0.904 0.967 0.984
Back2Color ms 192x640 0.092 0.685 4.229 0.170 0.911 0.968 0.984
Monodepth2[8] ms 320x1024 0.106 0.806 4.630 0.193 0.876 0.958 0.980
HR-Depth[25] ms 320x1024 0.101 0.716 4.395 0.179 0.894 0.966 0.983
CADepth[24] ms 320x1024 0.096 0.694 4.264 0.173 0.908 0.968 0.984
MonoViT[33] ms 320x1024 0.093 0.671 4.202 0.169 0.912 0.969 0.984
Back2Color ms 320x1024 0.087 0.644 4.079 0.166 0.920 0.970 0.985
@@
Method Sup. Error Metric Accuracy Metric
@@
Back2Color DIFFNet 0.094 0.660 4.254 0.172 0.904 0.967 0.984
@@
Back2Color VAN2 0.092 0.685 4.229 0.170 0.911 0.968 0.984
@@
Additionally, we used various synthetic datasets such as vKitti2, SHIFT, and UrbanSyn to real-world unsupervised depth model training, as shown in Tab.VII. The results consistently achieved excellent improvement, demonstrating the strong adaptability of our method. To demonstrate that the improvement in our Back2Color method is not due to the increased training samples (Synthetic Datasets) but rather due to the diverse priors brought by various synthetic datasets, we conducted experiments using 20% and 50% synthetic samples (Rows 6, 7, and Rows 9, 10) while ensuring the total number of dataset samples remains the same as the KITTI dataset (39,810). The results are consistent with those obtained using the full dataset, proving that the enhancement of our method is not due to the increase in training samples. Additionally, when using the MixSyn mixed synthetic dataset, some metrics (rmse_log and a1) show significant improvement compared to using a single synthetic dataset, demonstrating that our method benefits from learning the variety of multiple synthetic datasets.
Finally, to demonstrate the exceptional efficiency of our model, we compared it with the current best-performing MonoViT [33] and the lightweight Lite-mono [34]. Our VAN0 model is not only lighter than Lite-mono but also delivers superior performance. Additionally, it achieves competitive results with only one-sixth of the parameters of MonoViT. Meanwhile, our VAN2 model, which has a parameter count similar to MonoViT, shows a significant performance improvement, making it the current state-of-the-art (SOTA) model.
TABLE VII: Ablation Study on the impact of Various Synthetic Datasets.
Size Syn Number of Samples Error Metric Accuracy Metric
Kitti vKitti2 SHIFT UrbanSyn abs_rel sq_rel rmse rmse_log a1 a2 a3
192x640 vKitti2 39810 21260 0 0 0.094 0.694 4.296 0.173 0.907 0.968 0.984
192x640 SHIFT 39810 0 153000 0 0.092 0.676 4.238 0.171 0.909 0.969 0.984
192x640 Urban 39810 0 0 7539 0.093 0.644 4.198 0.172 0.907 0.968 0.984
192x640 MixSyn 39810 21260 153000 7539 0.092 0.685 4.229 0.170 0.911 0.968 0.984
320x1024 vKitti2 39810 21260 0 0 0.090 0.677 4.173 0.169 0.914 0.969 0.984
320x1024 vKitti2 31848 7962 0 0 0.090 0.675 4.143 0.168 0.917 0.970 0.985
320x1024 vKitti2 19905 19905 0 0 0.091 0.684 4.192 0.169 0.913 0.969 0.985
320x1024 SHIFT 39810 0 153000 0 0.089 0.651 4.158 0.167 0.916 0.970 0.985
320x1024 SHIFT 31848 0 7962 0 0.090 0.684 4.158 0.168 0.916 0.970 0.984
320x1024 SHIFT 19905 0 19905 0 0.090 0.680 4.144 0.167 0.915 0.970 0.985
320x1024 Urban 39810 0 0 7539 0.090 0.654 4.084 0.167 0.915 0.970 0.985
320x1024 MixSyn 39810 21260 153000 7539 0.087 0.644 4.079 0.166 0.920 0.970 0.985
@@ IV-E Visualization Analysis
TABLE VIII: Ablation study on models with params and computational complexities(GFLOPs).
@@
MonoViT small 320x1024 33.63 159.14 0.093 0.671 4.202 0.169 0.912 0.969 0.985
Lite-mono 8M 320x1024 8.76 29.88 0.097 0.710 4.309 0.174 0.905 0.967 0.984
Back2Color VAN0 320x1024 5.17 17.64 0.094 0.706 4.257 0.171 0.910 0.969 0.984
Back2Color VAN1 320x1024 18.61 51.48 0.090 0.664 4.139 0.168 0.915 0.970 0.984
Back2Color VAN2 320x1024 31.32 76.01 0.087 0.644 4.079 0.166 0.920 0.970 0.985
Back2Color VAN3 320x1024 49.5 114.63 0.087 0.653 4.068 0.165 0.919 0.971 0.985
Back2Color VAN0 384x1280 5.17 17.64 0.093 0.676 4.169 0.171 0.911 0.969 0.984
Back2Color VAN1 384x1280 18.61 51.48 0.090 0.664 4.102 0.168 0.917 0.970 0.984
Back2Color VAN2 384x1280 31.32 76.01 0.086 0.629 4.054 0.165 0.920 0.970 0.985
Back2Color VAN3 384x1280 49.5 114.63 0.087 0.652 4.062 0.165 0.920 0.970 0.984
@@ IV-F3 Can the VAN-based encoder replace the ResNet structure in the Pose encoder
so far, we have not observed any new research that can replace the ResNet-based PoseEncoder[8], whether using CNN-based or Transformer-based methods. We have attempted some alternatives, but none have improved depth estimation performance; in fact, many have resulted in performance degradation. In this study, we also tried using a VAN-based model to replace the ResNet-based PoseEncoder, but still did not observe any significant improvement. We believe that the pose estimation task is a small parameter regression task that is relatively simple compared to depth prediction and does not require a powerful network.